Consider the following equilibrium:


CH4 + H2O = CO + 3H2      Keq = 5.67 


An equilibrium mixture of this system was found to contain the following concentrations: 

[CH4] = 0.59 M, [H2O] = 0.63 M, [CO] = 0.25 M


What was the equilibrium [H2]?




1
Expert's answer
2021-06-28T08:10:53-0400

Solution:

Balanced chemical equation:

CH4 + H2O → CO + 3H2

The equilibrium expression for the reaction system:



Keq = 5.67

[CH4] = 0.59 M

[H2O] = 0.63 M

[CO] = 0.25 M

[H2] = unknown


Therefore,



[H2]3 = (5.67 × 0.59 × 0.63) / (0.25) = 8.43

[H2] = (8.43)1/3 = 2.035

[H2] = 2.035 M


Answer: [H2] = 2.035 M